OpenStudy (anonymous):

r=5cos(5theta) I need to find the area enclosed by one leaf. I know that this is a double integral problem, and I know that r goes from 0 to 5cos(5theta), but I don't know how to get the limits of integration for the angle...

OpenStudy (anonymous):

why a double integral?

OpenStudy (anonymous):

well I suppose it's the only way I know how to do it right now, but this is how the test question is gonna be. plus I don't know how to do regular integrals in polar anyway

OpenStudy (anonymous):

so the only way to do it is to have a picture of it?

OpenStudy (anonymous):

just take \[25\int_0^{\frac{\pi}{10}} \cos(5\theta) d\theta\]

OpenStudy (anonymous):

in general it is \[\int_a^b \frac{1}{2}r^2 d\theta\]

OpenStudy (anonymous):

i was integrating so that you go from 0 to pi/2, which gives you half the petal, then double. that takes care of the 1/2

OpenStudy (anonymous):

and of course pull the 5^2=25 out of the integral

OpenStudy (anonymous):

oh crap i had a typo. it should be cosine squared. sorry

OpenStudy (anonymous):

\[25\int_0^{\frac{\pi}{10}} \cos^2(5\theta) d\theta\]

OpenStudy (anonymous):

\[\int\limits_{\theta=a}^{\theta=b}\int\limits_{r=c}^{r=d}rdrd \theta\]

OpenStudy (anonymous):

have fun. but you do not need a double integral for this, that is for sure.

OpenStudy (anonymous):

is what it's supposed to look like, but I can't figure out the theta's limits. and 0 to pi/2 gives you 1.5 petals of the 5cos(5theta)

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

that is why i am integrating from 0 to pi/10

OpenStudy (anonymous):

you want \[5\theta=\frac{\pi}{2}\] ,making \[\theta=\frac{\pi}{10}\]

OpenStudy (anonymous):

so integrate from 0 to pi/10, not from 0 to pi/2

OpenStudy (anonymous):

why do you set the 5theta equal to pi/2?